OpenStudy (anonymous):

does x= 5 for 3^(2x+5)=27^x

OpenStudy (anonymous):

Correct! 27 = 3^3 Therefor: 3^(2x+5) = 3^3x ==> 2x+5 = 3x ==> x = 5

OpenStudy (cwrw238):

3^(2x+5) = 3^(3x) 2x + 5 = 3x x = 5